# Downsizing Diffusion Models for Cardinality Estimation

### v1

@@ 6.2.1. Accuracy Evaluation
The results for all estimators are listed in Table 1(a) for forest, power and Table 1(b) for modulo, which we shall now discuss.
@@
We begin by discussing the real-world datasets. On the most important benchmarks GM, 95th, and 99th Q-error, ADC+ and Naru consistently demonstrate top-tier performance, rivaling each other while significantly outperforming every other comparative technique. Meanwhile, ADC performs almost identically to ADC+ on 95th and 99th Q-error, but falls somewhat behind on GM Q-error due to its higher median Q-error values. As Naru, the only estimator which sometimes outperform ADC+, comes with the downside of much higher latency, such experimental results are enough to earn ADC+ a place among the best cardinality estimators.
Regarding max Q-error, the performance of all six estimators seem to be highly volatile and significantly influenced by random factors: for example, on forest, DeepDB performed notably better in our experiment than it did in (Wang et al., 2021), but its max Q-error increased from 1293 to 3856; on power, LW-NN performed roughly the same as in (Wang et al., 2021), but its max Q-error dropped from 40000 to 22000; meanwhile, while ADC’s largest Q-error was 551 on forest and 267 on power, it’s second-largest Q-error dropped sharply to 137 and 46.9, respectively, meaning that its max Q-error might change drastically on a slightly different workload. That said, ADC and ADC+ are still capable of outperforming every model except Naru and MSCN on max Q-error, proving ADC to be a relatively robust estimator.
@@
For the dataset modulo, ADC’s advantage scenario, the complex, multi-attribute correlations caused Naru, DEEPDB, and MSCN to significantly drop in accuracy, while LW-Tree’s performance exhibited less drastic declines. LW-NN is the only estimator to exhibit performance comparable to ADC on this dataset, almost catching up with ADC+ on 95th and 99th Q-error but still falling significantly behind ADC+ on the benchmarks GM and max. Meanwhile, the GMM predictor of ADC, a close relative of the KDE algorithm, only exhibits barely satisfactory performance on 95th and 99th Q-error, while Bayesnet implemented via the Chow-Liu algorithm, which only captures pairwise correlations, can be proven incapable of learning anything more than single-attribute histograms, showing that this performance decline is not unique to learned cardinality estimators.
Except for max Q-error on power, ADC consistently outperforms its standalone GMM predictor on GM, 95th, 99th, and max Q-error, but the magnitude of this improvement varies greatly across datasets: ADC reduces the distance of 99th Q-error to 1 by roughly one half on modulo, one fourth on forest, and one sixth on power. Thus, except for scenarios with intense latency requirements, using diffusion models for cardinality estimation will always pay off.
6.2.2. Latency and Storage Cost
@@
To evaluate the practicality of ADC+ in actual user scenarios, we discuss the inference latency and storage requirements of ADC+ in comparison with other learned models.
In terms of storage space, ADC+ is at an advantage compared to all other models, with parameters taking up 0.34MB of space for forest and less than 0.3MB for power and modulo. To be more precise, the network $s_{\theta_{1},head}$ usually takes up around 150 KB of storage, $s_{\theta_{2},tail}$ around 50 KB, the GMM predictor usually takes up around $10d$ KB with $d$ the dataset’s dimensionality, and other artifacts, including the decision tree, some database information, and a histogram we use for single-attribute queries usually take up around 20 KB. Meanwhile, all other models in our experiments have at least 0.42MB of trainable parameters (except Naru on power), and the actual number often goes further beyond that. We note that other learned models can be made more lightweight if needed, though often at a cost. For example, in their original essay, LW-Tree and LW-NN (Dutt et al., 2019) conducted their experiments using models as small as 16KB but, as a result, measured significantly higher Q-errors on forest and power than we did.
@@
Regarding per-query latency, the most important feature besides accuracy, ADC+ is much slower than MSCN, LW-Tree and LW-NN, which answer a query within 0.5 ms (within 0.2ms for LW-Tree), and somewhat slower than DeepDB, but is also $2\times$ as fast as Naru on forest and power while more than $4\times$ as fast as Naru on modulo. We find that ADC+’s slower speed on forest might be due to two factors: the heavier network we implemented due to the higher dimensionality and looser correlations, and the fact that ADC’s GMM predictor becomes less accurate as dimensionality rises, producing a satisfactory estimation by itself (hence allowing ADC+ to skip the sampling and correction phases) only on very rare occasions. Given that ADC+ is almost entirely written in Pytorch, it might also be possible to further speed up ADC+ with the help of GPUs which, sadly, we do not have access to.
@@
Figure 3. Performance of ADC and ADC’s GMM predictor with respect to GMM-estimated selectivity ($x$-axis) and query volume ($y$-axis), both shown on a log scale:
@@ 6.3. Reducing Median Q-error: Why ADC+
Our greatest motivation for creating ADC+ is that ADC, in its raw form, performs somewhat awkwardly on median Q-error, nearly always losing to its GMM predictor. In fact, on power and modulo, ADC ironically increased the distance between median Q-error and 1 by five times, ranking embarrassingly near bottom place on this benchmark despite its decent 95th, 99th, and max Q-error performance.
Why is this? We suspect that it’s caused by the fact that unlike other estimators, there are no ”very easy” queries for ADC, and even a query that only asks about two independent attributes or one whose queried region spans 90% of sample space must undergo the same prediction-sampling-correction procedure. In this case, ADC’s predictor-corrector integration scheme often finds itself adding non-negligible variance to correct an already tiny bias, a choice that is increasingly likely to backfire as the GMM predictor grows more competent.
To test our hypothesis, we create a scatterplot documenting the query volume and GMM-predicted selectivity of 100,000 training queries for each of the two real-world datasets, with blue points depicting queries where the GMM predictor performs better than ADC and red points depicting the opposite. Indeed, we can see from the experimental results documented in Figure 3 that while red points appear more frequently than blue in most regions of the graph, each graph’s upper-right corner always contains a very dense cluster that is almost exclusively blue.
What’s even better, separating a large proportion of blue points from red, at least those on the upper-right corner, can be done by drawing simple, clear-cut boundary lines: simple enough to be handled by a decision tree, which is exactly what ADC+ does. Compared to neural networks, a decision tree of depth 4 only takes up less than 3KB and classifies a query in less than 20 microseconds, allowing it to be implemented almost for free.
One might worry whether wrongly classifying points that actually benefit from predictor-corrector integration would lead to worse 95th and 99th percentile Q-error results. Fortunately, experiments show that ADC+’s effects on these benchmarks, though small, is almost never negative. On median Q-error, meanwhile, ADC+ acts as the game changer, successfully beating MSCN, LW-Tree, and LW-NN due to roughly halving the distance between median Q-error and 1 on all three datasets. What’s even better, ADC+ is also able to cut down the per-query latency of ADC by roughly 25%, because the benefits gained from potentially avoiding two lengthy steps of numerical integration often far outweighs the negligible classification costs.
@@ 7. Conclusion
The ability of diffusion models to reconstruct complex, high dimensional distributions makes them accurate cardinality estimators. We have proven this by constructing ADC, whose accuracy exceeds many of its famous peers on both real-world and synthetic datasets. Downsized using newfound theorems regarding the score function’s behavior, ADC also takes up less than 400KB of storage, and keeps its latency within 10ms on two out of three datasets tested.
ADC’s current capabilities are still somewhat limited by its inability to handle categorical attributes, an issue we seek to resolve in future studies. We are also curious whether some theorems, network structures, and modeling approaches we proposed upon designing ADC, such as switching from noise to data prediction as $t$ gets larger, might be beneficial for training diffusion models in general, despite influences of the manifold hypothesis and the much higher (3000+) dimensionality.
Overall, ADC provides a good starting point for exploring the use of diffusion models in cardinality estimation, a rather unconventional crossover that can produce very promising fruits.
Appendix A Mathematical Notations Used in the Essay
Table 2. Mathematical Notations Related to Diffusion Models
Notation Definition
$d$ Dimensionality of the dataset
$w$ The standard Brownian motion
$\hat{w}$ The standard Wiener process in reverse time
$p_{0}(x)$ The original data distribution we seek to reconstruct
$p_{t,\alpha,\beta}(x)$ The distribution derived by perturbing $p_{0}$ using the diffusion SDE $dx=-\alpha(t)dt+\beta(t)dw$ from time 0 to time $t$, also abbreviated as $p_{t}(x)$
$p_{0t,\alpha,\beta}(x|x_{0})$ The distribution derived by perturbing $\delta_{x_{0}}$, a distribution concentrated on the point $x_{0}$, using the above diffusion SDE from time 0 to time t, also abbreviated as $p_{0t}(x|x_{0})$.
$p_{t,appr,\alpha,\beta}(x)$ The distribution derived by perturbing $\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$, a distribution concentrated on the $N$ known data points, using the above SDE from time 0 to time $t$, also abbreviated as $p_{t,approx}(x)$
$q(x)$ Distribution reconstructed by the GMM predictor.
$q_{t,\alpha,\beta}(x)$ The distribution derived by perturbing $q_{0}=q(x)$ using the above diffusion SDE from time 0 to time t, also abbreviated as $q_{t}(x)$.
$\epsilon$ Early stopping time implemented upon training $s_{\theta}(x,t)$, see 3.3 for why early stopping is needed in score matching.
$\tilde{s}_{\theta}(x,t)$ $s_{\theta}(x,t+\epsilon)$
$k_{\alpha,\beta}(t)$ Function satisfying $e^{\int_{0}^{t}\alpha(s)ds}$, also abbreviated as $k(t)$
$\sigma^{2}_{\alpha,\beta}(t)$ Function satisfying the initial value ODE $\sigma^{2}(0)=0$, $\frac{d}{dt}(\sigma^{2}(t))=\beta(t)-2\alpha(t)\sigma^{2}(t)$, also abbreviated as $\sigma^{2}(t)$
$\varphi_{\sigma^{2}}$ Distribution density function of the 1-dimensional normal distribution $\mathcal{N}(0,\sigma^{2})$
$\tau_{k}$ The scaling operator $(\tau_{k}\circ g)(x)=k^{d}g(kx)$
$\Phi$ Cumulative density function of the 1-dimensional standard normal distribution $\mathcal{N}(0,1)$
Appendix B Proof of Theorems and Claims
Claim 1.
(Equivalence of perturbation schemes) Denote by $p_{0}(x)$ the underlying probability density function and $p_{t,\alpha,\beta}(x)$ the distribution of $x(t)$ when $x(0)\sim p_{0}(x)$ is perturbed according to the SDE
$dx=-\alpha(t)xdt+\beta(t)dw,t\in[0,T].$
Then, we would have
$p_{t,\alpha,\beta}$ $=(\tau_{k(t)}\circ p_{0})*\varphi_{\sigma^{2}(t)}$
$\nabla\log p_{t,\alpha,\beta}(x)$ $=k(t)\nabla\log p_{\sigma^{2}(t)k^{2}(t),0,1}(k(t)x),$
where $k(t)=e^{\int_{0}^{t}\alpha(s)ds}$, $\sigma(t)$ satisfies $\frac{d\sigma^{2}}{dt}=\beta(t)-2\alpha(t)\sigma^{2}(t)$ with initial condition $\sigma^{2}(0)=0$, and $\tau$ in $\mathbb{R}^{d}$ is the scaling operator $(\tau_{k}\circ f)(x)=k^{d}f(kx).$
@@ Appendix B Proof of Theorems and Claims
    $\forall x_{0}\in V_{i}$ such that $B(x,r)\subseteq V_{i}$, if $\nabla\log p_{i}(x)$ exists and is Lipschitz continuous in a neighborhood of $x_{0}$ with constant $h$, then
@@
    $\forall x_{0}\in\partial V_{i}\cap\partial V_{j}$ with at most one of $i$ and $j$ being zero, if $B(x_{0},r)\subseteq V_{i}\cup V_{j}$, $p_{0,i}(x_{0})+p_{0,j}(x_{0})>0$, and inside $B(x_{0},r)$,
@@
    $\forall x_{0}\in V_{0}$, suppose $y_{0}\in\bar{V_{i}}$ is the point closest to $x_{0}$ such that $y_{0}\in\cup_{i=1}^{k}\bar{V_{i}}$. If $P(x\in B_{r}(y)|x\sim p_{0})>M_{1}r^{k_{1}},\forall 0<r<R$ for some constants $M_{1},k_{1}$, while
@@
    with $\delta_{\sigma^{2}}(x_{0},t)$ a residual term of order $o(\frac{1}{\sigma(t)^{1+\epsilon}})$ for all $\epsilon>0$.
@@
Remark 0.
Our model, ADC, estimates the moment of transition from $O(\frac{1}{\sigma^{4}(t)k^{2}(t)})$ decay to $O(\frac{1}{\sigma^{6}(t)k^{4}(t)})$ decay using $\max\left\|x\right\|_{2}^{2}$ rather than $\mathrm{E}_{p_{0}(x)}\left\|x\right\|_{2}^{2}$. This is due to the fact that, for $p_{0}$ with non-bounded support, such a transition might be arbitrarily postponed even if $\mathrm{E}_{p_{0}(x)}\left\|x\right\|_{2}^{2}$ remains bounded.
@@
$\lim_{N\to\infty}\sigma^{4}(t)k^{2}(t)\mathrm{E}_{p_{t,N}(x)}\left\|\nabla\log p_{t,N}(x)+\frac{x}{\sigma^{2}(t)}\right\|_{2}^{2}=1$
@@
    The partition is dense enough that for the same $y$, fluctuations of $\Delta t_{i}\delta_{t,space}(x_{0},y)$ for different $t$ in $t_{i-1},t_{i}]$ is negligible compared to the total integration error
@@
    The error terms $\{\delta_{i,time}(x_{0})\}_{i=1}^{N}$ are negligible compared to $\{\delta_{\tilde{t}_{i},space}(x_{0},y(\tilde{t_{i}}))\}_{i=1}^{N}$

### v2

@@ 6.2.1. Accuracy Evaluation
The results for all estimators are listed in Table 2(a) for real-world datasets and Table 2(b) for modulo, which we shall separately discuss.
@@
For the real-world datasets, ADC+ consistently demonstrates top-tier performance, beating every model besides Naru on every benchmark except ”max” and ”median” on forest. Naru is the only model with performance rivaling ADC+’s, ranking above ADC+ on four benchmarks but falling slightly behind on nine others. Meanwhile, ADC performs almost identically to ADC+ on 95th and 99th Q-error, but falls somewhat behind on GM Q-error due to its higher median Q-error values. Considering that the estimator Naru comes with the downside of much higher latency, such results are enough to earn ADC+ a place among the best cardinality estimators.
@@
For the dataset modulo, ADC’s advantage scenario, the complex, multi-attribute correlations caused Naru, DeepDB, and MSCN to significantly drop in accuracy, while LW-Tree’s performance exhibited less drastic declines. LW-NN is the only estimator to exhibit performance comparable to ADC on this dataset, almost catching up with ADC+ on 95th and 99th Q-error, but still falling significantly behind ADC+ on the benchmarks GM and max. Meanwhile, the GMM predictor of ADC, a close relative of the KDE algorithm, only exhibits barely satisfactory performance on 95th and 99th Q-error, while Bayesnet implemented via the Chow-Liu algorithm, which only captures pairwise correlations, would be incapable of learning anything more than single-attribute histograms, showing that this decline is not unique to learned estimators.
Compared to its standalone GMM predictor, ADC’s diffusion model components result in significant accuracy gains for challenging queries and workloads, but generally fail to further enhance accuracy if the workload and/or query is already simple. That is, ADC significantly reduced max Q-error on all datasets, notably reduced 99th percentile Q-error on all datasets except for higgs, which is the most simple; reduced 95th percentile Q-error for the hardest datasets forest and modulo, but increased median Q-error on all four datasets tested. The reasons why this is the case, as well as what to do about this phenomenon, shall be discussed in subsection 6.3.
6.2.2. Latency, Storage Space, and Training Time
@@
To evaluate the practicality of ADC+ in actual user scenarios, we discuss the inference latency, training time, and storage requirements of ADC+ in comparison with other learned models.
@@
Regarding per-query latency, the most important feature besides accuracy, ADC+ usually answers a query within the 7-10 milliseconds range on all datasets except forest. This is much slower than MSCN, LW-Tree and LW-NN, which answer a query within 0.5 ms (within 0.2ms for LW-Tree), and somewhat slower than DeepDB, but is also $2\times$ as fast as Naru on forest and power while more than $4\times$ as fast as Naru on higgs and modulo. Given that ADC+ is almost entirely written in Pytorch, it might also be possible to further speed up ADC+ with the help of GPUs, which, sadly, we do not have access to.
In terms of storage space, ADC+ is at an advantage compared to all other models, with parameters taking up 0.44MB of space for forest and less than 0.35MB for all other datasets. To be more precise, the network $s_{\theta_{1},head}$ usually takes up 125 to 200 KB of storage, $s_{\theta_{2},tail}$ around 50 KB, the GMM predictor usually takes up around $10d$ KB with $d$ the dataset’s dimensionality, the Bayesnet component, if used, usually takes up around 40KB, and other artifacts, including the decision tree, some database information, and a histogram we use for single-attribute queries, usually take up around 20 KB. Meanwhile, all comparative techniques have at least 0.42MB of trainable parameters (except Naru on power), and the actual number often goes further beyond that. We note that other learned models can be made more lightweight if needed, though often at a cost. For example, in their original essay, LW-Tree and LW-NN (refeighteen) conducted their experiments using models as small as 16KB, but, as a result, measured significantly higher Q-errors on forest, power, and higgs than we did.
ADC is rather at a disadvantage when it comes to training, with its total training time amounting to 10 hours and only forcibly exiting from training then due to meeting our time constraint. While Naru and LW-NN can also take several hours to train, which is especially true for large datasets, training usually takes as little as 20 minutes for MSCN, 10 minutes for DeepDB, and is completed within one minute for LW-Tree. However, we also point out that the training process of ADC, which heavily relies on Pytorch, can be greatly sped up with the help of a GPU (which we don’t have access to), and we are currently working on enhanced score matching methods to cut down training time while producing just as accurate results.
@@
Figure 3. Performance of ADC and ADC- with respect to GMM-estimated selectivity ($x$-axis) and query volume ($y$-axis), both shown on a log scale:
@@ 6.3. Reducing Median Q-error: Why ADC+
Our motivation for creating ADC+ is that ADC, in its raw form, performs somewhat awkwardly on median Q-Error, nearly always losing to its GMM predictor, ADC-. In fact, on modulo, ADC ironically increased the distance between median Q-error and 1 by five times, ranking embarrassingly near bottom place on this benchmark despite its decent 95th, 99th, and max Q-error performance.
Why is this? We suspect that it’s caused by the fact that, unlike other estimators, there are no ”very easy” queries for ADC, and even a query that only asks about two independent attributes or one whose queried region spans 90% of the sample space must undergo the same prediction-sampling-correction procedure. In this case, ADC’s predictor-corrector integration scheme often finds itself adding non-negligible variance to correct an already tiny bias, a choice that is increasingly likely to backfire as the GMM predictor grows more competent.
To test our hypothesis, we create a scatterplot documenting the query volume and GMM-predicted selectivity of 100,000 training queries for each of the 4 datasets, with blue points depicting queries where the GMM predictor performs better than ADC, red points depicting the opposite, and green depicting queries where the two methods are tied (often due to both predicting a cardinality smaller than 1). Indeed, we can see from Figure 3 that while the relative proportion of red points tends to increase notably in the lower-left corner, each graph’s upper-right corner contains a very dense cluster that is almost exclusively blue.
What’s even better, separating a large proportion of blue points from red, at least those on the upper-right corner, can often be done by drawing simple, straight boundary lines, a task ADC performs by training a decision tree. Compared to neural networks, a decision tree of depth 4 takes up less than 3KB and classifies a query in 20 microseconds, making its implementation almost free of costs.
One might worry whether wrongly classifying points that actually benefit from predictor-corrector integration would lead to worse 95th and 99th percentile Q-error results. Fortunately, experiments show that ADC+’s effects on these benchmarks, though small, are seldom negative. On median Q-error, meanwhile, ADC+ acts as the game changer, successfully halving the distance between median Q-error and 1 on all four datasets while successfully beating MSCN, LW-Tree, and LW-NN on power, higgs, and modulo. What’s even better, ADC+ is also able to cut down the per-query latency of ADC, because the benefits gained from potentially avoiding two lengthy steps of numerical integration often far outweigh the negligible classification costs.
@@ 7. Conclusion
The ability of diffusion models to reconstruct complex distributions with high dimensions makes them accurate cardinality estimators. We have proven this by constructing ADC and ADC+, whose accuracy exceeds many famous peers on real-world and synthetic datasets.
The current capabilities of ADC+ are slightly limited by its long training time and inability to handle categorical attributes, issues we seek to resolve in future studies. We also wonder whether some theorems, network structures, and modeling approaches we proposed upon designing ADC, such as switching from noise to data prediction as $t$ gets larger, might be beneficial for training diffusion models in general, despite influences of the manifold hypothesis and the much higher (3000+) dimensionality.
Overall, ADC+ provides a good starting point for exploring the use of diffusion models in cardinality estimation, a rather unconventional crossover that can produce very promising fruits.
Appendix A Proof of Theorems
Theorem 2.1 0.
Let $p_{0}$ be a probabilistic distribution and $\{p_{t}|t\in[0,T]\}$ be a family of distributions derived by perturbing $p_{0}$ using the SDE (2) from time 0 to time $T$. Let $k(t)=e^{\int_{0}^{t}\alpha(s)ds}$, and $\sigma^{2}(t)$ be the solution of the initial value Ordinary Differential Equation (ODE)
${d\sigma^{2}}/{dt}$ $=\beta(t)-2\alpha(t)\sigma^{2}(t)$
$\sigma^{2}(0)$ $=0,$
then, for each choice of $\alpha$ and $\beta$, it would hold that
(12) $p_{t,\alpha,\beta}$ $=(\tau_{k(t)}\circ p_{0})*\varphi_{\sigma^{2}(t)}$
where $\tau$ in $\mathbb{R}^{d}$ is the scaling operator $(\tau_{k}\circ f)(x)=k^{d}f(kx)$ and $\varphi_{\sigma^{2}(t)}$ in $\mathbb{R}^{d}$ is the probability density function of $\mathcal{N}(0,\sigma^{2}(t)I_{d})$.
@@ Appendix A Proof of Theorems
Let $s_{\theta}(x,t):\mathbb{R}^{d}\times\mathbb{R^{+}}\to\mathbb{R}^{d}$ be a neural network and $p_{0}$ be a probability distribution, and $p_{0,\theta}(x)$ to be the marginal distribution of $\tilde{x}(0)$ after $\tilde{x}(T)\sim\mathcal{N}(0,\sigma^{2}(T)I_{d})$ is perturbed by the reverse diffusion SDE
$dx=-[\alpha(t)x+\beta^{2}(t)s_{\theta}(x,t)]dt+\beta(t)d\hat{w}$
from time $t$ to time 0. Using the notation of Theorem Theorem 2.1, the KL divergence between $p_{0}$ and $p_{\theta}$ can be upper bounded by
while the value of $\log p_{\theta}(x_{0})$ can be lower bounded by
$\log p_{0,\theta}(x_{0})\geq\mathrm{E}_{p_{0T}(x|x_{0})}\log\varphi_{\sigma^{2}(T)}(x)+\int_{0}^{T}n\alpha(t)dt-\frac{1}{2}\int_{0}^{T}\beta^{2}(t)\cdot$
(15) $\mathrm{E}_{p_{0t(x|x_{0})}}$ $[\left\|\nabla\log p_{0t}(x|x_{0})-s_{\theta}(x,t)\right\|_{2}^{2}-\left\|\nabla\log p_{0t}(x|x_{0})\right\|_{2}^{2}]dt$
with the equal sign holding in both inequalities if there exists a probability distribution $q$ such that $s_{\theta}(x,t)=\nabla\log q_{t}(x),\forall t\in[0,T]$.
Proof.
Proof for this theorem is elegantly provided by Song, Durkan, Murray and Ermon in (reftwo). ∎
Theorem 3.1 0.
@@
    $\forall x\in V_{i}$ such that $B(x,r)\subseteq V_{i}$,
@@
    with $\delta_{\sigma^{2}}(x_{0},t)$ a term of order $o(\frac{1}{\sigma(t)^{1+\epsilon}})$ for all $\epsilon>0$.
@@
Note: Our model, ADC, estimates the moment of transition from $O(\frac{1}{\sigma^{4}(t)k^{2}(t)})$ decay to $O(\frac{1}{\sigma^{6}(t)k^{4}(t)})$ decay using $\max\left\|x\right\|_{2}^{2}$ rather than $\mathrm{E}_{p_{0}(x)}\left\|x\right\|_{2}^{2}$. This is due to the fact that, for $p_{0}$ with non-bounded support, such a transition might be arbitrarily postponed even if $\mathrm{E}_{p_{0}(x)}\left\|x\right\|_{2}^{2}$ remains bounded.
@@
$\lim_{N\to\infty}\sigma^{4}(t)k^{2}(t)\mathrm{E}_{p_{t,N}(x)}\left\|\nabla\log p_{t,N}(x)+\frac{x}{\sigma^{2}(t)}\right\|_{2}^{2}=1.$
@@
    The partition is dense enough that for the same $y$, fluctuations of $\Delta t_{i}\delta_{t,space}(x_{0},y)$ for different $t$ in $t_{i-1},t_{i}]$ is negligible compared to the total integration error,
@@
    The error terms $\{\delta_{i,time}(x_{0})\}_{i=1}^{N}$ are negligible compared to $\{\delta_{\tilde{t}_{i},space}(x_{0},y(\tilde{t_{i}}))\}_{i=1}^{N}$, and